OpenStudy (meilendurcer):

i'm really bad at this kind of math please help. If g(x) = x2 + 3, find g(4).

OpenStudy (anonymous):

what u do is plug in 4 as x so it will look like this g(4)=(4)^2+3 g(4)=8+3 g(4)=11

OpenStudy (nirmalnema):

replace x with 4 =(4)^2+3 16+3 g(4)=19

OpenStudy (anonymous):

G(x) is the same thing as putting a y there, but we put g(x) so that it is easier to see what we put in for x, well at least that's what i was taught.

OpenStudy (anonymous):

oh yea the answer should have been 19 i messed up on the 4^2

OpenStudy (meilendurcer):

i was kind of wondering why there was a smaller number than what the equation looks like. thanks.

OpenStudy (nirmalnema):

your welcome